# Aligned Novel View Image and Geometry Synthesis via Cross-modal Attention Instillation

### v2

@@ Abstract
We introduce a diffusion-based framework that performs aligned novel view image and geometry generation via a warping-and-inpainting methodology. Unlike prior methods that require dense posed images or pose-embedded generative models limited to in-domain views, our method leverages off-the-shelf geometry predictors to predict partial geometries viewed from reference images, and formulates novel-view synthesis as an inpainting task for both image and geometry. To ensure accurate alignment between generated images and geometry, we propose cross-modal attention distillation (MoAI) where attention maps from the image diffusion branch are injected into a parallel geometry diffusion branch during both training and inference. This multi-task approach achieves synergistic effects, facilitating geometrically robust image synthesis as well as well-defined geometry prediction. We further introduce proximity-based mesh conditioning to integrate depth and normal cues, interpolating between point cloud and filtering erroneously predicted geometry from influencing the generation process. Empirically, our method achieves high-fidelity extrapolative view synthesis on both image and geometry across a range of unseen scenes, delivers competitive reconstruction quality under interpolation settings, and produces geometrically aligned colored point clouds for comprehensive 3D completion. Project page is available at https://cvlab-kaist.github.io/MoAI/.
Figure 1: Our diffusion-based framework generates novel view image and its geometry at arbitrary target poses, from single or multiple unposed input reference images. With an off-the-shelf geometry module, we predict an partial point cloud whose RGB values and geometry are projected to target viewpoint. Our diffusion models inpaint from these partial observations to generate plausible novel view image and geometry, with cross-Modal Attention Instillation (MoAI) ensuring alignment between modalities, combined with partial geometry resulting a complete 3D scene.
@@ 1 Introduction
Novel view synthesis (NVS), the task of reconstructing 3D scenes from sparse 2D reference images, represents a fundamental challenge that requires neural networks to understand and model the underlying 3D structure of the world from limited 2D observations. Seminal works such as NeRF [20] and 3DGS [14] implicitly model volumetric geometry and appearance by optimizing scene-specific representations to fit individual 3D scenes from reference images. To generalize beyond single-scene optimization, feedforward methods [29, 3, 4, 31] have emerged for direct 3D prediction. In addition, the advent of diffusion models [24] has introduced generative NVS methods [6, 25, 23] that achieve remarkable fidelity in novel view synthesis.
@@
However, numerous challenges persist. While feedforward methods [29, 3, 4, 31] demonstrate high fidelity in interpolative settings by accurately reconstructing regions visible in reference images, they lack extrapolation capabilities - the ability to synthesize occluded or unseen regions. On the other hand, generative NVS methods [6, 25, 23], trained to generate novel views from predetermined camera pose embeddings, are capable of extrapolation. However, these approaches are limited in synthesizing novel views at arbitrary out-of-domain viewpoints [27] that are underrepresented in the training pose embedding distribution. Consequently, these methods also require known reference camera poses as input, limiting them to posed image novel view synthesis settings.
In this context, we build upon warping-and-inpainting methods [5, 25] to introduce an alternative approach to multi-view novel view synthesis. These methods leverage off-the-shelf geometry prediction models [13, 29, 28] to estimate geometry and camera pose from a reference image, project this geometry to a given target viewpoint, and use this projection to guide the generative process for plausible target view synthesis.
@@
We propose to predict geometry from multiple reference views, aggregating and projecting them to a novel viewpoint as coarse geometric conditioning that guides spatial cross-attention mechanisms within diffusion networks. By interpreting NVS as an inpainting problem in this manner, our approach enables synthesis at arbitrary viewpoints from unposed reference images, with reconstruction and generative capabilities at extrapolative viewpoints. Additionally, we extend this method to novel view geometry synthesis by training a geometry denoising U-Net that similarly inpaints geometry from reference view geometries. This approach offers a significant advantage: whereas previous depth prediction methods struggle with scale-shift discrepancy [2, 13] — inconsistency of depth scale and offset between predicted and ground truth geometry — our method achieves geometric alignment with reference geometry by generating as a continuation of the reference geometry.
@@
However, we observe that the image inpainting network struggles to synthesize large regions with no reference projections, while the geometry completion network handles such regions robustly—our analysis demonstrates that this disparity stems from the more deterministic nature of geometry completion compared to image inpainting. To counter this issue, we introduce cross-modal attention instillation, shortened as MoAI, in which spatial attention maps from the image denoising network — which implicitly model cross-view correspondences — are transferred to replace the attention maps within the geometry network, ensuring cross-modal alignment. This creates a synergistic multi-task learning framework: the inherently deterministic geometry completion regularizes image generation, while the geometry network leverages rich semantic features from images for improved synthesis quality. Additionally, our proximity-based mesh conditioning incorporates additional geometric cues (depth and normal) into correspondence conditions, facilitating sparse geometry interpolation and filtering of erroneous warped conditions.
@@
Our method demonstrates powerful extrapolation capabilities for both novel view image and geometry synthesis, resulting in aligned colored point clouds that achieve 3D scene completion (Fig. 1). Our results show it achieves state-of-the-art performance in extrapolative camera settings while maintaining competitive reconstruction performance and zero-shot generalization to unseen data.
@@ 2.1 Non-generative few-shot NVS methods
Optimization methods. Neural 3D representations such as NeRF [20] and 3DGS [14] require numerous calibrated views to optimize the neural radiance field effectively. Optimization-based few-shot methods [10, 21, 15, 16] alleviate this issue by tailoring a single 3D scene from sparse views. For instance, DietNeRF [10] enforces semantic consistency between rendered images from novel viewpoints and available reference images, while RegNeRF [21] regularizes the geometry and appearance of patches from unobserved viewpoints. However, these approaches suffer from limited generalization ability beyond individual scenes and the computational burden of per-scene optimization.
@@
Feedforward methods. Feedforward NVS approaches [32, 4, 29, 31, 8] address few-shot novel view synthesis without per-scene optimization. PixelNeRF [32] is among the first to condition a NeRF on image inputs using local CNN features, predicting novel view image in a feedforward manner. MVSplat [4] improves on this by predicting 3D Gaussians from sparse multi-view images with a cost volume for depth estimation, yielding high-quality 3D representations. Subsequent works [29, 17, 8, 31] tackle the pose-free scenario, where networks predict novel views from unposed images; for example, DUSt3R [29] and MASt3R [17] leverage transformers to output the pointmaps and estimate camera poses, and Noposplat [31] jointly predicts 3D representations and poses from sparse inputs. However, these methods generally lack extrapolative capabilities, unable to synthesize novel geometry or appearance in regions unseen or occluded from the reference images.
2.2 Few-shot NVS with generative models
@@
Recent diffusion-based NVS approaches [18, 27, 26, 6] leverage their generative capacity to synthesize novel views from one or few images. Zero-123 [18] fine-tunes a diffusion model to directly generate a novel viewpoint image for a relative pose from a single image. Similarly, MVDream [26] and CAT3D [6] employ spatial cross-attention between the generating viewpoints to achieve consistent novel view synthesis at target viewpoints. Although these diffusion-based approaches offer strong extrapolative capability, they do not provide explicit geometry for the novel viewpoints, necessitating a separate optimization process on NeRF or 3DGS for full geometry. Moreover, because they receive target view camera pose as a feature embedding, the range of poses they can generate is limited to the training domain, hindering the direct generation of arbitrary novel poses.
2.3 Warping-and-inpainting methods
@@
Diffusion models [24] have demonstrated strong inpainting capabilities [19], which has motivated their application to novel view synthesis (NVS). LucidDreamer [5] leverages off-the-shelf monocular depth estimators [2, 29, 28] to extract geometry from a single image, warp it to a target viewpoint, and inpaint missing regions, while GenWarp [25] uses predicted geometry as a correspondence signal for implicit inpainting. However, because these methods rely primarily on 2D image inpainting without a comprehensive understanding of 3D structure, they struggle to synthesize scenes with large view differences, particularly in object-centric scenarios where novel view results in warped geometry covering only a small portion of the target viewpoints.
@@ 3 Method
Given $N$ unposed and sparse reference RGB images $\{I_{n}\in\mathbb{R}^{H\times W\times 3}\}_{n=1}^{N}$, with height $H$ and weight $W$, the method’s objective is the joint prediction of novel view image $I_{t}$ and pointmap $P_{t}$ for target viewpoint $\pi_{t}$, leveraging the diffusion model’s generative capabilities for high-fidelity novel view and geometry synthesis. Our method extends warping-and-inpainting methodology [17, 25] from single-image to multi-view settings. We then also apply the same approaches to the geometry domain, performing geometry completion at the target viewpoint from partial geometry predicted by off-the-shelf models (Sec. 3.1). To ensure alignment between target image and geometry at target view, we introduce cross-modal attention distillation (Sec. 3.2), a multitask learning that yields synergistic benefits for both modalities. Finally, to handle noise and artifacts in predicted point clouds, we introduce proximity-based mesh conditioning (Sec. 3.3), which prevents erroneous artifacts from degrading generation quality.
@@ 3.1 Novel view image generation
Our image generation architecture, as shown in the upper section of Fig. 2, consists of two U-Nets: an image reference network and an image denoising network. The image reference network [9, 25] extracts semantic reference features from the reference images $\{I_{n}\in\mathbb{R}^{H\times W\times 3}\}_{n=1}^{N}$, and the image denoising network utilizes the features from the reference network to generate a novel-view image.
Geometry prediction and pointmap projection. We first leverage an off-the-shelf geometry prediction model [29, 28] to obtain the set of corresponding camera poses $\mathcal{\varphi}=\{\pi_{n}\in\mathbb{R}^{4\times 4}\}_{n=1}^{N}$ as well as pointmap $\{P_{n}\in\mathbb{R}^{H\times W\times 3}\}_{n=1}^{N}$ from reference images. The pointmap $P_{n}$ [29] is a 2D grid of 3D point coordinates, where each element represents the predicted world coordinate for the given pixel. Next, the pointmaps of the reference images, $P_{1},P_{2}\dots P_{N}$, are interpreted as unordered sets of 3D points then merged into a single point cloud $P$. We then project $P$ onto the target viewpoint $\pi_{t}$:
@@
resulting in the projected pointmap $P^{\Pi}_{t}$ for the target view $\pi_{t}$, where $\Pi(\cdot)$ denotes a projection function. When multiple points project onto a single pixel, only the point closest to the target image plane is rendered, as in the standard point cloud rasterization procedure [25].
Pointmap correspondence conditioning. We leverage the projected pointmap $P^{\Pi}_{t}$ and the reference view pointmaps $P_{n}$ as a sparse geometric correspondence condition, which enables the image denoising network to implicitly establish the correspondences between the target viewpoint and reference images. Specifically, we first encode $P^{\Pi}_{t}$ using positional embedding $\mathcal{E}(\cdot)$ and concatenate the resulting Fourier feature $\mathcal{E}(P^{\Pi}_{t})$ with a binary mask $M_{t}$, which marks the grid pixels without any projected points [25]. This forms the target correspondence condition $\mathbf{c}^{\mathbf{t}}$ for image denoising network at target viewpoint $\pi_{t}$. Similarly, for each reference viewpoint $\pi_{n}$, we obtain a reference correspondence condition $\mathbf{c}^{\mathbf{r}}_{n}$ which consists of an embedded reference view pointmap Fourier feature $\mathcal{E}(P_{n})$, concatenated with a one-valued tensor mask $\mathbf{1}$, since every grid pixel in the reference view pointmap has a corresponding 3D point due to dense prediction by the off-the-shelf model and therefore marked as 1. As we obtain a reference correspondence condition for every reference image $\{I_{n}\}_{n=1}^{N}$:
@@
Similarly to Hu et al. [9], these correspondence conditions are first passed through a convolutional network, resulting in target and reference correspondence condition features. The target correspondence condition feature is added to the target image latent feature from the first convolutional layer of the image denoising network, while each reference correspondence condition feature is similarly added to the features of its corresponding reference image within the image reference network. Such conditioning guides the image denoising network in identifying relevant spatial correspondences from multiple reference images to ensure consistency in novel view generation. Notice that instead of providing explicit pixel-to-pixel correlation (e.g., warped pixel coordinates [25]) between reference viewpoints and target viewpoint, we directly provide embedded pointmap as condition. This design choice allows the model to associate each spatial location in the target image with multiple potential correspondences, providing the reference images for robust reconstruction.
Aggregated attention. We conduct an aggregated attention between the target view image features derived from the image denoising network and the reference features produced by the image reference network. We acquire the image key features $K^{I}_{t}\in\mathbb{R}^{1\times C\times(W\times H)}$ and image value features $V^{I}_{t}\in\mathbb{R}^{1\times C\times(W\times H)}$ from the spatial self-attention layers of the image denoising network. These target view image key and value features are concatenated with $N$ image key and value reference features, resulting in combined image key feature $K^{I}$ and combined image value feature $V^{I}$. Then we conduct aggregated attention [9, 25] with $K^{I}$, $V^{I}$, and the image target view feature $Q^{I}_{t}$ as a query:

### v3

@@ Abstract
We introduce a diffusion-based framework that generates aligned novel view images and geometries via a warping‐and‐inpainting methodology. Unlike prior methods that require dense posed images or pose-embedded generative models limited to in‐domain views, our method leverages off‐the‐shelf geometry predictors to predict partial geometries viewed from reference images, and formulates novel view synthesis as an inpainting task for both image and geometry. To ensure accurate alignment between the generated image and geometry, we propose cross-Modal Attention Instillation (MoAI) where the attention maps from an image diffusion branch are injected into a parallel geometry diffusion branch during both training and inference. This multi-task approach achieves synergistic effects, facilitating both geometrically robust image synthesis and geometry prediction. We further introduce proximity‐based mesh conditioning to reduce erroneous projections and to integrate depth and normal cues to the correspondence conditions. Empirically, our method achieves high-fidelity extrapolative view synthesis, delivers competitive reconstruction under interpolation settings, and produces geometrically aligned point clouds as 3D completion. Code and weights will be publicly released.
Figure 1: Overview of our diffusion-based framework. From one or more unposed reference images, we predict a partial colored point cloud and project it to the target view. Our diffusion model then inpaints missing regions with the cross-Modal Attention Instillation (MoAI), ensuring alignment between image and geometry, resulting in a complete 3D scene.
@@ 1 Introduction
Novel view synthesis (NVS), the task of reconstructing 3D scenes from sparse 2D reference images, represents a fundamental challenge that requires neural networks to understand and model the underlying 3D structure of a scene from limited 2D observations. Seminal works such as NeRF (Mildenhall et al., 2021) and 3DGS (Kerbl et al., 2023) implicitly model volumetric geometry and appearance by optimizing scene-specific representations to fit individual 3D scenes from reference images. To generalize beyond single-scene optimization, feedforward methods (Wang et al., 2024; Charatan et al., 2024; Chen et al., 2024; Ye et al., 2024) have emerged for direct 3D prediction. In addition, the advent of diffusion models (Rombach et al., 2022) has introduced generative NVS methods (Gao et al., 2024; Seo et al., 2024; Ren et al., 2025) that achieve remarkable fidelity in NVS.
@@
However, significant challenges remain. Feedforward methods (Wang et al., 2024; Charatan et al., 2024; Chen et al., 2024; Ye et al., 2024) show high fidelity in interpolative settings by filling the regions visible in reference images, but they lack extrapolation capabilities for synthesizing occluded or unseen areas. Conversely, generative NVS methods (Gao et al., 2024; Seo et al., 2024; Ren et al., 2025), trained with known camera poses, can extrapolate beyond the reference views. However, when conditioned on camera viewpoints underrepresented during training, these models often produce erroneous novel views (Voleti et al., 2024). Consequently, these methods require known reference camera poses, limiting them to the posed NVS setting.
@@
Building on warping-and-inpainting methods (Chung et al., 2023; Seo et al., 2024), we propose an alternative framework for multi-view NVS. This framework leverages an off-the-shelf geometry prediction model (Ke et al., 2024; Wang et al., 2024; 2025) to estimate geometry and camera pose from reference images, project this geometry to a target viewpoint, and guide the generative process for plausible NVS. More specifically, we predict geometry from multiple reference views, aggregate and project them to a novel viewpoint, and use this coarse geometric conditioning to guide spatial cross-attention in diffusion networks. By framing NVS as an inpainting problem, our approach synthesizes novel views at arbitrary viewpoints from unposed reference images, while also supporting reconstruction and generation at extrapolative viewpoints. Additionally, we extend this method to novel view geometry synthesis by training a geometry denoising U-Net that inpaints the target geometry from reference views. This offers a key advantage: unlike prior depth-prediction methods Chung et al. (2023); Seo et al. (2024) that suffer from a scale-shift discrepancy between predicted and known reference geometry (Ke et al., 2024), our method ensures geometric alignment by generating as a continuation of the reference geometry.
@@
To encourage synergistic multi-task learning across image and depth modalities, ensuring their alignment, we introduce cross-Modal Attention Instillation, shortened as MoAI, where spatial attention maps from the image denoising network, which implicitly capture cross-view correspondences, are instilled to replace those of the geometry network during both training and inference. More specifically, our design enables to learn from a relatively robust and consistent geometry completion task to regularize image generation through an instilled attention map, while the geometry network leverages rich semantic features from images to improve synthesis quality. Additionally, our proximity-based mesh conditioning incorporates additional geometric cues (depth and normal) into correspondence conditions, interpolating sparse geometry and filtering of erroneous projections.
@@
Our method demonstrates strong extrapolation capabilities for both novel view image and geometry synthesis, resulting in aligned colored point clouds that achieve 3D scene completion (Fig.˜1). It achieves state-of-the-art performance in extrapolative settings, while maintaining competitive reconstruction and zero-shot generalization to unseen data.
@@ 2 Related work
Neural 3D representations such as NeRF (Mildenhall et al., 2021) and 3DGS (Kerbl et al., 2023) require numerous calibrated views to optimize the neural radiance field effectively. Optimization-based few-shot methods (Jain et al., 2021; Niemeyer et al., 2022; Kim et al., 2022; Kwak et al., 2023) alleviate this issue by tailoring a single 3D scene from sparse views. For instance, DietNeRF (Jain et al., 2021) enforces semantic consistency between rendered images from novel viewpoints and available reference images, while RegNeRF (Niemeyer et al., 2022) regularizes the geometry and appearance of patches from unobserved viewpoints. However, these approaches are unable to generalize beyond individual scenes and are computationally expensive.
@@
Feedforward NVS approaches (Yu et al., 2021; Chen et al., 2024; Wang et al., 2024; Ye et al., 2024; Hong et al., 2023) address few-shot novel view synthesis without per-scene optimization. PixelNeRF (Yu et al., 2021) is among the first to condition a NeRF on image inputs using local CNN features, predicting a novel view image in a feedforward manner. MVSplat (Chen et al., 2024) improves on this by predicting 3D Gaussians from sparse multi-view images with a cost volume for depth estimation, yielding high-quality 3D representations. Subsequent works (Wang et al., 2024; Leroy et al., 2024; Hong et al., 2023; Ye et al., 2024) tackle the pose-free scenario, where networks predict novel views from unposed images; for example, DUSt3R (Wang et al., 2024) and MASt3R (Leroy et al., 2024) leverage transformers to output the pointmaps and estimate camera poses, and Noposplat (Ye et al., 2024) jointly predicts 3D representations and poses from sparse inputs. However, these methods generally lack extrapolative capabilities, unable to synthesize novel geometry or appearance in regions unseen or occluded from the reference images.
Generative few-shot NVS.
@@
Recent diffusion-based NVS approaches (Liu et al., 2023; Voleti et al., 2024; Shi et al., 2023; Gao et al., 2024) leverage their generative capacity to synthesize novel views from one or a few images. Zero123 (Liu et al., 2023) fine-tunes a diffusion model to directly generate a novel viewpoint image for a relative pose from a single image, and ZeroNVS (Sargent et al., 2023) extends upon this for single-view novel view synthesis. Similarly, MVDream (Shi et al., 2023) and CAT3D (Gao et al., 2024) employ spatial cross-attention between the generating viewpoints to achieve consistent novel view synthesis at target viewpoints, while ViewCrafter (Yu et al., 2024) delivers high-fidelity performance by finetuning a video diffusion model. Although these approaches offer strong extrapolative capability, they do not provide explicit geometry for the novel viewpoints, necessitating a separate optimization process on NeRF or 3DGS for full geometry. Moreover, because they receive target view camera pose as a feature embedding, the range of poses they can generate is limited to the training domain, hindering the direct generation of arbitrary novel poses.
Warping-and-inpainting.
@@
Diffusion models (Rombach et al., 2022) have demonstrated strong inpainting capabilities (Lugmayr et al., 2022), which has motivated their application to novel view synthesis (NVS). LucidDreamer (Chung et al., 2023) leverages off-the-shelf monocular depth estimators (Bhat et al., 2023; Wang et al., 2024; 2025) to extract geometry from a single image, warp it to a target viewpoint, and inpaint missing regions, while GenWarp (Seo et al., 2024) uses predicted geometry as a correspondence signal for implicit inpainting. However, because these methods rely primarily on 2D image inpainting without a comprehensive understanding of 3D structure, they struggle to synthesize scenes with large view differences, particularly in object-centric scenarios where novel view results in warped geometry covering only a small portion of the target viewpoints.
@@ 3 Method
Given $N$ unposed and sparse reference RGB images $\{I_{n}\in\mathbb{R}^{H\times W\times 3}\}_{n=1}^{N}$, with height $H$ and weight $W$, the method’s objective is the joint prediction of novel view image $I_{t}$ and pointmap $P_{t}$ for target viewpoint $\pi_{t}$, leveraging the diffusion model’s generative capabilities for high-fidelity novel view and geometry synthesis. Our method extends the warping-and-inpainting methodology (Leroy et al., 2024; Seo et al., 2024) from single-image to multi-view settings. Importantly, we generalize this strategy from the image domain to geometry, performing geometry completion at the target viewpoint from partial geometry predicted by off-the-shelf models (Sec. 3.1). To ensure alignment between the target image and geometry, we introduce cross-modal attention distillation (Sec. 3.2), a multitask learning that yields synergistic benefits for both modalities. Finally, to handle noise and artifacts in predicted point clouds, we introduce proximity-based mesh conditioning (Sec. 3.3), which prevents erroneous artifacts from degrading generation quality.
@@ 3.1 Novel view image generation
Our image generation architecture, as shown in the upper section of Fig. 2, consists of two U-Nets: an image reference network and an image denoising network. The image reference network (Hu et al., 2023) extracts semantic reference features from the reference images $\{I_{n}\in\mathbb{R}^{H\times W\times 3}\}_{n=1}^{N}$, and the denoising network utilizes the features from the reference network to generate a novel view image.
Geometry prediction and pointmap projection. We first leverage an off-the-shelf geometry prediction model (Wang et al., 2024; 2025) to obtain the set of corresponding camera poses $\mathcal{\varphi}=\{\pi_{n}\in\mathbb{R}^{4\times 4}\}_{n=1}^{N}$ as well as pointmap $\{P_{n}\in\mathbb{R}^{H\times W\times 3}\}_{n=1}^{N}$ from reference images. The pointmap $P_{n}$ (Wang et al., 2024) is a 2D grid of 3D point coordinates, where each element represents the predicted world coordinate for the given pixel. Next, the pointmaps of the reference images, $P_{1},P_{2}\dots P_{N}$, are interpreted as unordered sets of 3D points then merged into a single point cloud $P$. We then project $P$ onto the target viewpoint $\pi_{t}$:
@@
resulting in the projected pointmap $P^{\Pi}_{t}$ for the target view $\pi_{t}$, where $\Pi(\cdot)$ denotes a projection function. When multiple points project onto a single pixel, only the point closest to the target image plane is rendered, as in the standard point cloud rasterization procedure (Seo et al., 2024).
Pointmap correspondence conditioning. We leverage the projected pointmap $P^{\Pi}_{t}$ and the reference view pointmaps $\{P_{n}\}^{N}_{n=1}$ as a sparse geometric correspondence condition, which enables the image denoising network to establish the correspondences between the target viewpoint and reference images. Specifically, we first encode $P^{\Pi}_{t}$ using positional embedding $\mathcal{E}(\cdot)$ and concatenate the resulting Fourier feature $\mathcal{E}(P^{\Pi}_{t})$ with a binary mask $M_{t}$, which marks the grid pixels without any projected points. This forms the target correspondence condition $\mathbf{c}^{\mathbf{t}}$ for the image denoising network at target viewpoint $\pi_{t}$. Similarly, for each reference viewpoint $\pi_{n}$, we obtain a reference correspondence condition $\mathbf{c}^{\mathbf{r}}_{n}$ which consists of an embedded reference view pointmap Fourier feature $\mathcal{E}(P_{n})$, concatenated with a one-valued tensor mask $\mathbf{1}$, since every grid pixel in the reference view pointmap has a corresponding 3D point due to dense prediction by the off-the-shelf model and therefore marked as 1. As we obtain a reference correspondence condition for every reference image $\{I_{n}\}_{n=1}^{N}$:
@@
Similar to Hu et al. (2023), these correspondence conditions are first passed through a convolutional network, resulting in target and reference correspondence condition features. The target correspondence condition feature is added to the target image latent feature from the first convolutional layer of the image denoising network, while each reference correspondence condition feature is similarly added to the features of its corresponding reference image within the image reference network. Such conditioning guides the image denoising network in identifying relevant spatial correspondences from multiple reference images to ensure consistency in novel view generation. Notice that instead of providing explicit pixel-to-pixel correlation (e.g., warped pixel coordinates (Seo et al., 2024)) between reference viewpoints and target viewpoint, we directly provide an embedded pointmap as a condition. This design choice allows the model to associate each spatial location in the target image with multiple potential correspondences, providing the reference images for robust reconstruction.
Aggregated attention. We conduct an aggregated attention between the target view image features derived from the image denoising network and the reference features produced by the image reference network. We acquire the image key features $K^{I}_{t}\in\mathbb{R}^{1\times C\times(W\times H)}$ and image value features $V^{I}_{t}\in\mathbb{R}^{1\times C\times(W\times H)}$ from the spatial self-attention layers of the image denoising network. These target view image key and value features are concatenated with $N$ image key and value reference features, resulting in the combined image key and value features, $K^{I}$ and $V^{I}$, respectively. Then we conduct aggregated attention (Seo et al., 2024) with $K^{I}$, $V^{I}$, and the image target view feature $Q^{I}_{t}$ as a query: